# Polynomial Composition Activations: Unleashing the Dynamics of Large Language Models

### v1

@@ 5 Related Work
Polynomial activation functions (Hornik et al., 1989; Oh et al., 2003), although less commonly used, have been studied in various contexts for their ability to model higher-order, complex relationships more effectively. For instance, Lokhande et al. (2020) introduced Hermite polynomial activations to improve pseudo-label accuracy, while Chrysos et al. (2020) proposed polynomial networks, $\Pi$-nets, which apply to various domains such as image and audio processing. Building on this, Chrysos et al. (2023) utilized regularization techniques to enhance the performance of polynomial networks. These works highlight the potential of polynomial functions to increase the expressiveness of neural networks by capturing intricate, higher-order interactions. On the theoretical front, the expressivity and approximation power of polynomial functions have been rigorously explored (Kileel et al., 2019; Kidger & Lyons, 2020; Kubjas et al., 2024).
@@ A.1 Proof of Lemma 2
First, we observe that $\mathrm{PolyReLU}(x)=\mathrm{Poly}(\mathrm{ReLU}(x)$, where $Poly(x)=\sum_{i=0}^{r}a_{i}x^{i}$ for $x\in[-1,1]$. By Lemma A.1, there exists a ReLU network $f_{1}:[0,1]\rightarrow[-1,1]$ of size $O(\ln^{2}(1/\epsilon))$ such that:
@@
Thus, we construct $f=f_{1}\circ\mathrm{ReLU}$ for inputs $x\in[-1,1]$. This yields that:
@@
A.2 Proof of Theorem 1
The proof is an elementary extension of Lemma 1.
Proof of Theorem 1.
A.3 Proof of Theorem 2
@@
Denote $g_{i}$ as the $i$-th layer of PolyReLU neteeork $f$ for $1\leq i\leq L$, such that:
@@
For each neuron, since $\|a\|_{1}+b\leq 1$, it follows that:
@@
Additionally, note that the range of PolyReLU is $[-1,1]$. Hence, by induction, the output of each neuron remains within $[-1,1]$. For each subnetwork $g_{i}$, by applying Lemma 2, we can construct a corresponding ReLU network $f_{i}$ by replacing each PolyReLU activation $p_{i,j}$ in $g_{i}$ with a ReLU activation. Specifically, for any $i\in[L]$ [^4] We use notation $[L]$ to denote the set $\{1,2,\dots,L\}$. and $j\in[K]$, there exists a ReLU network $f_{i,j}:[-1,1]\rightarrow[-1,1]$ that approximates the PolyReLU activation $p_{i,j}$ with given tolerance $\epsilon_{i}>0$.
Thus, the network $f_{i}$ is obtained by replacing each PolyReLU activation $p_{i,j}$ in $g_{i}$ with its ReLU approximation $f_{i,j}$. Obviously, $f_{i}$ is a ReLU network whose output dimensions are in range $[-1,1]$.
@@
Therefore, we derive the following approximation bound:
@@
for $\forall i\in[L]$. Since $h^{g}_{0}=h^{f}_{0}$, we have $e_{0}=0$. Let $\epsilon_{i}=\epsilon/(L\alpha^{L-i})$ for $\forall i\in[L]$. It follows that:
@@
Last, we need to estimate the size of the ReLU network $f$. By Lemma 2, the size of each ReLU subnetwork $f_{i,j}$ is $O(\ln^{2}(L\alpha^{L-i}/\epsilon))$. Therefore, the total size of the ReLU network $f$ is:
@@
where we use the fact that:
@@ A.4 Proof of Theorem 3
Let $M\in{\mathbb{N}}$ and $\rho:{\mathbb{R}}\rightarrow{\mathbb{R}}$ be any continuous piece-wise linear function with $M$ breakpoints. Then there exists a PolyReLU network $g$ of size $O(M)$ such that:
@@
By expressing $n$ in base $r$, we have that:
@@
$g(x)=x^{n}\quad\forall x\in{\mathbb{R}}.$
@@
where ${\bm{m}}=(m_{1},m_{2},\dots,m_{d})\in\{-N,-(N-1),\dots,0,\dots,N\}$, and $\varphi$ is defined as:
@@
The approximation error between $f$ and $f_{N}$ can be bounded as follows:
@@
Next, we construct a PolyReLU network $g_{N}$ to represent $f_{N}$ exactly. Let $a_{{\bm{m}},{\bm{n}}}=\frac{1}{{\bm{n}}!}D^{{\bm{n}}}f\left(\frac{{\bm{m}}}{N}\right)$. Since $\|f\|_{{\mathcal{W}}^{n,\infty}\left([-1,1]^{d}\right)}\leq 1$, $|a_{{\bm{m}},{\bm{n}}}|\leq 1$ for any ${\bm{m}},{\bm{n}}$, we rewrite $f_{N}$ as:
@@
Therefore, $f_{N}$ is composed of at most $d^{n}(2N+1)^{d}$ functions $\phi_{{\bm{m}}}({\bm{x}})\left({\bm{x}}-\frac{{\bm{m}}}{N}\right)^{{\bm{n}}}$. Since $\phi_{{\bm{m}}}({\bm{x}})=\prod_{i=1}^{d}\varphi\left(3N\left(x_{k}-\frac{m_{k}}{N}\right)\right)$ and each $\varphi\left(3N\left(x_{k}-\frac{m_{k}}{N}\right)\right)$ is a continuous piece-wise linear function, we can apply Corollary A.1, which guarantees that there exists a PolyReLU network $\hat{\phi}_{{\bm{m}}}$ of size $O(d)$ that can exactly represent $\phi_{{\bm{m}}}$ on ${\mathbb{R}}^{d}$, i.e., $\hat{\phi}_{{\bm{m}}}({\bm{x}})=\phi_{{\bm{m}}}({\bm{x}}),\forall{\bm{x}}\in{\mathbb{R}}^{d}$. For $\left({\bm{x}}-\frac{{\bm{m}}}{N}\right)^{{\bm{n}}}=\prod_{i=1}^{d}\left(x_{i}-\frac{m_{i}}{N}\right)^{n_{i}}$, by Lemma A.4, we know that there exists a PolyReLU network $g_{{\bm{m}}}$ of size at most $O(d\ln^{2}(n))$ such that $g_{{\bm{m}}}({\bm{x}})=\left({\bm{x}}-\frac{{\bm{m}}}{N}\right)^{{\bm{n}}},\forall{\bm{x}}{\mathbb{R}}^{d}$. Combining these results, we can now construct a larger PolyReLU network $g_{n}$ as follows:
@@
where the total size of the network is:
@@
Here, we use Eq. (30) to determine the size bound in terms of the error tolerance $\epsilon$. Clearly, we have:
@@
Appendix B Activation Functions
@@
SiLU(Swish) SiLU$(x)=x*\sigma(x)$
(Ramachandran et al., 2017)
@@
Poly $\mathrm{Poly}(x)=\sum_{i=0}^{r}a_{i}x^{i}$, $a_{i},i\in[r]$ are learnable parameters.
Appendix C Experimental Details
C.1 Architecture
@@
Hidden Layers Exports Context Length
16 64 4096
C.2 Definition of Effective Rank
We adopt the concept of effective rank from Roy & Vetterli (2007) to measure the effective dimensionality of a matrix. Given a matrix $A$ with Singular Value Decomposition (SVD) $A=U\Sigma V^{\top}$, where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{n}\geq 0$. we define the singular value distribution as $p_{i}=\sigma_{i}/\sum_{j=0}^{n}\sigma_{j},i\in[n]$. The effective rank of $A$ is then given by:
@@ C.2 Definition of Effective Rank
Appendix D Additional Results on Dense Model

### v2

@@ 5 Related Work
Polynomial activation functions (Hornik et al., 1989; Oh et al., 2003), although less commonly used, have been studied in various contexts for their ability to model higher-order, complex relationships more effectively. For instance, Lokhande et al. (2020) introduced Hermite polynomial activations to improve pseudo-label accuracy, while Chrysos et al. (2020) proposed polynomial networks, $\Pi$-nets, which apply to various domains such as image and audio processing. Building on this, Chrysos et al. (2023) utilized regularization techniques to enhance the performance of polynomial networks. These works highlight the potential of polynomial functions to increase the expressiveness of neural networks by capturing intricate, higher-order interactions. On the theoretical front, the expressivity and approximation power of polynomial functions have been rigorously explored (Kileel et al., 2019; Kidger & Lyons, 2020; Kubjas et al., 2024). Additionally, (Li et al., 2019) investigated the approximation capabilities of rectified power units (i.e., ReLU2), demonstrating that they achieve the same approximation rate as PolyReLU.
@@ 6 Conclusions
Acknowledgments
Jinwen Ma was supported by the Natural Science Foundation of China under grant 62071171.
@@ A.1 Proof of Lemma 1
Proof of Lemma 1.
Thus, for any polynomial activation of order $r$,
A.2 Proof of Theorem 1
The proof is an elementary extension of Lemma 1.
Proof of Theorem 1.
A.3 Proof of Lemma 2
@@
First, we observe that $\mathrm{PolyReLU}(x)=\mathrm{Poly}(\mathrm{ReLU}(x)$, where $Poly(x)=\sum_{i=0}^{r}a_{i}x^{i}$ for $x\in[-1,1]$. By Lemma A.1, there exists a ReLU network $f_{1}:[0,1]\rightarrow[-1,1]$ of size $O(\ln^{2}(1/\epsilon))$ such that
@@
Thus, we construct $f=f_{1}\circ\mathrm{ReLU}$ for inputs $x\in[-1,1]$. This yields that
@@
A.4 Proof of Theorem 2
@@
Denote $g_{i}$ as the $i$-th layer of PolyReLU neteeork $f$ for $1\leq i\leq L$, such that
@@
For each neuron, since $\|a\|_{1}+b\leq 1$, it follows that
@@
Additionally, note that the range of PolyReLU is $[-1,1]$. Hence, by induction, the output of each neuron remains within $[-1,1]$. For each subnetwork $g_{i}$, by applying Lemma 2, we can construct a corresponding ReLU network $f_{i}$ by replacing each PolyReLU activation $p_{i,j}$ in $g_{i}$ with a ReLU activation. Specifically, for any $i\in[L]$ [^4] We use the notation $[L]$ to denote the set $\{1,2,\dots,L\}$. and $j\in[K]$, there exists a ReLU network $f_{i,j}:[-1,1]\rightarrow[-1,1]$ that approximates the PolyReLU activation $p_{i,j}$ with given tolerance $\epsilon_{i}>0$.
Thus, the network $f_{i}$ is obtained by replacing each PolyReLU activation $p_{i,j}$ in $g_{i}$ with its ReLU approximation $f_{i,j}$. Obviously, $f_{i}$ is a ReLU network whose output dimensions are in the range $[-1,1]$.
@@
Therefore, we derive the following approximation bound
@@
for $\forall i\in[L]$. Since $h^{g}_{0}=h^{f}_{0}$, we have $e_{0}=0$. Let $\epsilon_{i}=\epsilon/(L\alpha^{L-i})$ for $\forall i\in[L]$. It follows that
@@
Last, we need to estimate the size of the ReLU network $f$. By Lemma 2, the size of each ReLU subnetwork $f_{i,j}$ is $O(\ln^{2}(L\alpha^{L-i}/\epsilon))$. Therefore, the total size of the ReLU network $f$ is
@@
where we use the fact that
@@ A.5 Proof of Theorem 3
Let $M\in{\mathbb{N}}$ and $\rho:{\mathbb{R}}\rightarrow{\mathbb{R}}$ be any continuous piece-wise linear function with $M$ breakpoints. Then there exists a PolyReLU network $g$ of size $O(M)$ such that
@@
By expressing $n$ in base $r$, we have that
@@
$g(x)=x^{n},\quad\forall x\in{\mathbb{R}}.$
@@
where ${\bm{m}}=(m_{1},m_{2},\dots,m_{d})\in\{-N,-(N-1),\dots,0,\dots,N\}^{d}$, and $\varphi$ is defined as
@@
The approximation error between $f$ and $f_{N}$ can be bounded as follows
@@
Next, we construct a PolyReLU network $g_{N}$ to represent $f_{N}$ exactly. Let $a_{{\bm{m}},{\bm{n}}}=\frac{1}{{\bm{n}}!}D^{{\bm{n}}}f\left(\frac{{\bm{m}}}{N}\right)$. Since $\|f\|_{{\mathcal{W}}^{n,\infty}\left([-1,1]^{d}\right)}\leq 1$, $|a_{{\bm{m}},{\bm{n}}}|\leq 1$ for any ${\bm{m}},{\bm{n}}$, we rewrite $f_{N}$ as
@@
Therefore, $f_{N}$ is composed of at most $d^{n}(2N+1)^{d}$ functions $\phi_{{\bm{m}}}({\bm{x}})\left({\bm{x}}-\frac{{\bm{m}}}{N}\right)^{{\bm{n}}}$. Since $\phi_{{\bm{m}}}({\bm{x}})=\prod_{i=1}^{d}\varphi\left(3N\left(x_{k}-\frac{m_{k}}{N}\right)\right)$ and each $\varphi\left(3N\left(x_{k}-\frac{m_{k}}{N}\right)\right)$ is a continuous piece-wise linear function, we can apply Corollary A.1, which guarantees that there exists a PolyReLU network $\hat{\phi}_{{\bm{m}}}$ of size $O(d)$ that can exactly represent $\phi_{{\bm{m}}}$ on ${\mathbb{R}}^{d}$, i.e., $\hat{\phi}_{{\bm{m}}}({\bm{x}})=\phi_{{\bm{m}}}({\bm{x}}),\forall{\bm{x}}\in{\mathbb{R}}^{d}$. For $\left({\bm{x}}-\frac{{\bm{m}}}{N}\right)^{{\bm{n}}}=\prod_{i=1}^{d}\left(x_{i}-\frac{m_{i}}{N}\right)^{n_{i}}$, by Lemma A.4, we know that there exists a PolyReLU network $g_{{\bm{m}}}$ of size at most $O(d\ln^{2}(n))$ such that $g_{{\bm{m}}}({\bm{x}})=\left({\bm{x}}-\frac{{\bm{m}}}{N}\right)^{{\bm{n}}},\forall{\bm{x}}\in{\mathbb{R}}^{d}$. Combining these results, we can now construct a larger PolyReLU network $g_{n}$ as follows
@@
where the total size of the network is
@@
Here, we use Eq. (30) to determine the size bound in terms of the error tolerance $\epsilon$. Clearly, we have
@@
Appendix B Discussion of the Optimal Approximation Rate
For convenience, we state Theorem 4.2 in DeVore et al. (1989) in the following.
Theorem 4 (Theorem 4.2 in DeVore et al. (1989)).
Let ${\mathcal{X}}$ be a Banach space $L_{q}$ on ${\mathbb{R}}^{d}$, $1\leq q\leq\infty$. If $F_{n,d}^{p}=\{f\in{\mathcal{X}}|\|f\|_{{\mathcal{W}}^{n,p}}\leq 1\},1\leq p\leq q,n\in{\mathbb{N}}$, then
$\sup_{f\in F_{n,d}^{p}}\inf_{\theta\in{\mathbb{R}}^{m}}\|f-{\mathcal{M}}(\theta)\|_{q}\geq Cm^{-\frac{n}{d}},$ (36)
where ${\mathcal{M}}$ be a mapping from ${\mathbb{R}}^{m}$ into ${\mathcal{X}}$ which associate with each $\theta\in{\mathbb{R}}^{m}$ the element ${\mathcal{M}}(\theta)\in{\mathcal{X}}$, and $C$ is a constant.
Particularly, let $q=p=\infty$ and ${\mathcal{X}}=L_{\infty}[-1,-1]^{d}$, the above theorem tells us that the approximation error of the neural networks with $m$ parameters to approximate $F_{n,d}^{\infty}$, i.e., $F_{d,n}$, is larger than $Cm^{-\frac{n}{d}}$. Therefore, given error tolerance $\epsilon$, we have
$\epsilon\geq Cm^{-\frac{n}{d}}.$ (37)
It follows that
$m\geq C^{\frac{d}{n}}\epsilon^{-\frac{d}{n}}.$ (38)
Hence, the total number of parameters required by neural networks to approximate functions in $F_{n,d}$ is $\Omega(\epsilon^{-\frac{d}{n}})$. Combining with Theorem 3, we have that our PolyReLU networks achieve the optimal approximation rate in the context of Sobolev spaces.
Appendix C Activation Functions
@@
SiLU(Swish) (Ramachandran et al., 2017) SiLU$(x)=x*\sigma(x)$
@@
Poly $\mathrm{Poly}(x)=\sum_{i=0}^{r}a_{i}x^{i}$, $a_{i},i\in[r]$ are learnable parameters
Appendix D PyTorch Implementation of PolyCom
PyTorch implementations of PolyReLU and PolyNorm are provided in the following.
Algorithm 1 PyTorch-Style Implementation of PolyReLU
import torch
from torch.utils.checkpoint import checkpoint
import torch.nn.functional as F
def _poly(x, weight, bias, order=3):
return sum(weight[i] * (x ** (i+1)) for i in range(order)) + bias
class PolyReLU(torch.nn.Module):
def __init__(self):
super(PolyReLU, self).__init__()
self.weight = torch.nn.Parameter(torch.ones(3) / 3)
self.bias = torch.nn.Parameter(torch.zeros(1))
def forward(self, x, checkpointing=True):
x = F.relu(x)
if checkpointing:
return checkpoint(_poly, x, self.weight, self.bias, use_reentrant=False)
return _poly(x, self.weight, self.bias)
Algorithm 2 PyTorch-Style Implementation of PolyNorm
def _norm(x, eps=1e-6):
return x * torch.rsqrt(x.pow(2).mean(-1, keepdim=True) + eps)
def _poly_norm(x, weight, bias, order=3):
return sum(weight[i] * _norm(x ** (i+1)) for i in range(order)) + bias
class PolyNorm(torch.nn.Module):
def __init__(self):
super(PolyNorm, self).__init__()
self.weight = torch.nn.Parameter(torch.ones(3) / 3)
self.bias = torch.nn.Parameter(torch.zeros(1))
def forward(self, x, checkpointing=True):
if checkpointing:
return checkpoint(_poly_norm, x, self.weight, self.bias, use_reentrant=False)
return _poly_norm(x, self.weight, self.bias)
Appendix E Experimental Details
E.1 Architecture
@@
Hidden Layers Exports Context Length Weight Tying
16 64 4096 no
E.2 Hyperparameters
In Table 7, we list the hyperparameters that we use by default at training time for all our experiments for the 1B dense model and MoE-1B-7B, unless stated otherwise.
Table 7: Pretraining hyperparameters for the 1B dense model and MoE-1B-7B.
1B dense model MoE-1B-7B
Optimizer AdamW AdamW
Learning Rate (LR) 3E-4 4E-4
Minimum LR 3E-5 5E-5
LR Schedule cosine cosine
Weight Decay 0.1 0.1
$\beta_{1}$ 0.9 0.9
$\beta_{2}$ 0.95 0.95
Gradient Clipping 1 1
Warmup Tokens 620,$0$
Warmup Steps - 2000
Init Distribution normal trunc normal
Init std $1/\sqrt{2.5d}$ $1/\sqrt{2.5d}$
Init Truncation - 3$\times$ std
Load Balancing Loss Weight - 0.01
Router z-loss Weight - 0.001
E.3 Definition of Effective Rank
We adopt the concept of effective rank from Roy & Vetterli (2007) to measure the effective dimensionality of a matrix. Given a matrix $A$ with Singular Value Decomposition (SVD) $A=U\Sigma V^{\top}$, where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{n}\geq 0$. we define the singular value distribution as $p_{i}=\sigma_{i}/\sum_{j=0}^{n}\sigma_{j},i\in[n]$. The effective rank of $A$ is then given by
@@ E.3 Definition of Effective Rank
Appendix F Computational complexity analysis
For the sake of simplicity, we only calculate the computational complexity in one-layer Feed-Forward Networks (FNN) since activation only. Support input tensor of FFN is $x\in\mathbb{R}^{B\times S\times H}$, where $B$, $L$, and $H$ are the batch size, length of the sequence, and hidden size, respectively. Roughly, the relationship between computational FLOPs and model parameters can be regarded as proportional  [^5] https://blog.eleuther.ai/transformer-math/. Therefore, we can estimate the proportion of the computational cost incurred by the activation function calculations within the total computational cost of the FFN matrix computations ($24BSH^{2}$). The FLOPs ratio is calculated as
FLOPs ratio=FLOPs for $\text{FLOPs ratio}=\frac{\text{FLOPs for activation}}{24BSH^{2}}.$
It is important to note that the overhead and proportion often vary for different model sizes, so we provide the corresponding formulas directly and take $H=1024$, $B=4$ (each device), $S=4096$, using BF16 precision as an example. For PolyReLU and PolyNorm, we use the 3-order default setting. The results are summarized in the following tables:
Table 8: Comparison of computational complexity for different activation functions without gradient checkpointing.
Method Intermediate Size FLOPs for activation FLOPs ratio Memory Overhead
ReLU $4H$ $4BSH$ $\frac{1}{6H}=0.016\%$ $4BSH=128MB$
GeLU $4H$ $72BSH$ $\frac{3}{H}=0.29\%$ $10BSH=320MB$
SwiGLU $\frac{8}{3}H$ $\frac{112}{3}BSH$ $\frac{14}{9H}=0.15\%$ $8BSH=256MB$
ReLU2 $4H$ $8BSH$ $\frac{1}{3H}=0.032\%$ $8BSH=256MB$
PolyNorm $4H$ $72BSH$ $\frac{3}{H}=0.29\%$ $12BSH=384MB$
PolyReLU $4H$ $40BSH$ $\frac{5}{3H}=0.16\%$ $8BSH=256MB$
Table 9: Comparison of computational complexity for different activation functions with gradient checkpointing.
Method Intermediate Size FLOPs for activation FLOPs ratio Memory Overhead
ReLU $4H$ $8BSH$ $\frac{1}{3H}=0.033\%$ 0
GeLU $4H$ $144BSH$ $\frac{6}{H}=0.59\%$ 0
SwiGLU $\frac{8}{3}H$ $\frac{224}{3}BSH$ $\frac{28}{9H}=0.30\%$ 0
ReLU2 $4H$ $16BSH$ $\frac{2}{3H}=0.065\%$ 0
PolyNorm $4H$ $144BSH$ $\frac{6}{H}=0.59\%$ 0
PolyReLU $4H$ $80BSH$ $\frac{10}{3H}=0.33\%$ 0
We assume that the scale of the input tensor is set to $[-1,1]$. In this case, the FLOPs for both tanh and exp are approximately 10 each. For a fair comparison, the intermediate size of models with SwiGLU activations is set to 8/3H to keep the overall numbers of parameters constant.
In practice, we utilized gradient checkpointing  [^6] https://pytorch.org/docs/stable/checkpoint.html to reduce the additional memory overhead to 0. While this may introduce a certain computational overhead, given the overall modest computational cost of the activation functions, the overall increase in GPU memory and computational cost is quite small.
Appendix G Scaling Curves
Figure 8: Scaling curves of models with different activation functions.
In Figure 8, we present the training loss scaling curves for dense models utilizing the activation functions SwiGLU, PolyReLU, and PolyNorm. As illustrated in the figure, both PolyReLU and PolyNorm consistently outperform SwiGLU across model sizes ranging from 110M to 1.3B parameters.
The model sizes used for the scaling law experiments are detailed in Table 10, and all models employ the hyperparameters specified for 1B dense models, as listed in Table 7. Models with 110M, 226M, and 502M parameters were trained on 200B tokens.
Table 10: Model sizes for scaling laws experiments.
110M 768 2048 2048/3072 16 12
226M 1024 2048 2560/3840 16 16
502M 1536 2048 4096/6144 16 16
1.3B 2048 4096 5504/8256 16 24
Appendix H Experiments on Vision
To evaluate the effectiveness of PolyCom beyond language modeling, we trained ResNet50 (He et al., 2016) on ImageNet-1K (Deng et al., 2009) following the settings of timm (Wightman, 2019). For comparison, we replaced the ReLU activation in ResNet50 with PolyNorm and reported the training loss and top-1/top-5 accuracy on the evaluation set, as shown in Figure 9. The results demonstrate that PolyNorm outperforms ReLU by a significant margin in terms of training loss, top-1 accuracy, and top-5 accuracy. Specifically, PolyNorm achieves a lower training loss of 2.026 compared to ReLU’s 2.121, and improves top-1 and top-5 accuracy to 75.117% and 92.099%, surpassing ReLU by +0.204 and +0.068, respectively.
Figure 9: Training loss and evaluation accuracy on ImageNet-1K for ResNet50 models with ReLU and PolyNorm activations. PolyNorm achieves lower training loss and higher top-1/top-5 accuracy, demonstrating improved performance.
Appendix I Additional Results on Dense Model